OpenStudy (anonymous):

Solve the system, using substitution. What is the x coordinate of the solution? x = 2y - 2 and 3x - y = 14 A. 6 B. 4 C. 1 D. 0 E. no solution

OpenStudy (nirmalnema):

x = 2y - 2 y=3x-14 put the value of in the above equation x=2(3x-14)-2 x=6x-28-2 -5x=-30 x=6

OpenStudy (nirmalnema):

@Abiilovee911 do u get that

OpenStudy (anonymous):

yes thank you (: !

OpenStudy (nirmalnema):

welcome..!!